OpenStudy (anonymous):

If the sin 30° = , then which statement is true? Select one: a. cos 150° = 0 because the cosine and sine are supplements b. cos 60° = because the cosine and sine are complements c. cos 150° = 1 because the cosine and sine are supplements d. cos 60° = 0 because the cosine and sine are complements

OpenStudy (anonymous):

Sin 30 deg=?

OpenStudy (anonymous):

sin 30 = 1/2

OpenStudy (whpalmer4):

Probably the problem showed 1/2 as a character which did not copy/paste, leaving a blank spot...

OpenStudy (anonymous):

Assuming that you take standard value of sin(30) which is 1/2 answer would be b because all the other options cannot be true. sin and cos are complements. That is cos(90-theta)=sin(theta) or vice-versa. To make this simple remember that the angles of sin and cos should together make 90 deg.

OpenStudy (anonymous):

cosine 60 = 1/2

OpenStudy (anonymous):

b. is the answer @savvvvannahh

OpenStudy (anonymous):

Yes where it said sin 30 degrees = 1/2 was supposed to be there. I copied it heehe :3 But B is still the answer? @sayakdbz @silversurfer

OpenStudy (anonymous):

yes indeed @savvvvannahh